OpenStudy (anonymous):

Help please! Which of the following equations has 3 as a solution? -2b=6 4a-5=12 5y=y+8 z-15=-12 Thank you!!

jimthompson5910 (jim_thompson5910):

Let's say we want to check 4a-5=12 So plug in a = 3 (replace 'a' with 3) to get... 4a-5=12 4*3-5=12 12 - 5 = 12 7 = 12 ... which is FALSE, so a = 3 is NOT the solution to 4a-5=12

OpenStudy (anonymous):

z-15=-12